THE VIBRATIONAL SPECTRUM OF THE DIMER OF FORMIC ACID IN THE GAS PHASE: ROOM TEMPERATURE AND JET–COOLED FT–DATA

R. GEORGES, Equipe Astrochimie Expérimentale, Laboratoire PALMS, UMR CNRS 6627, Université de Rennes I, Campus de Beaulieu, F-35042 Rennes, France; M. FREYTES,^{a} D. HURT-MANS^{b}, Service de Chimie Quantique et Photophysique C. P. 160/09, Université Libre de Bruxelles, 50 Avenue F. D. Roosevelt, B–1050 Brussels, Belgium; I. KLEINER, Laboratoire de Photophysique Moléculaire, CNRS, Université Paris-Sud, Campus d’Orsay, Bât. 350, F–91405 Orsay Cedex, France; J. VANDER AUWERA, \underline{M. HERMAN}, Service de Chimie Quantique et Photophysique C. P. 160/09, Université Libre de Bruxelles, 50 Avenue F. D. Roosevelt, B–1050 Brussels, Belgium.

the jet.

bA.R.C. Researcher, Belgium

Absorption spectra of the gaseous (C_{2h}) dimer of formic acid were recorded from 20 to 11300 cm^{−1} using a high resolution Fourier transform spectrometer, at room temperature. Jet-cooled spectra were recorded over the range 1850 to 3750 cm^{−1} and liquid phase data were recorded in the NIR region. Some of the observed bands are vibrationally assigned, significantly extending the information available in the literature. The rotational structure of the ν_{15} fundamental c–type band in the FIR is partly unraveled and its hot band pattern considered in more details. Band contour simulation is performed for the jet-cooled ν_{6} + ν_{21} band, in the MIR, allowing the rotational temperature to be estimated to 35K, in

aPresent address: CEILAP, Centro de Investigaciones en Láseres y Aplicaciones (CITEFA, CONICET), Zufriategui 4380, 1603-Villa Martelli, Buenos Aires, Argentina